OpenStudy (anonymous):

Solve for x. 4x^2 -4x = 3 A. {3,-4.} B. {-3/2,4} C. {-1/2,4} D.{3/2,-1/2} My answer: D show me how I got this answer

OpenStudy (anonymous):

well wat do u think tell me and i will help u kk

OpenStudy (anonymous):

I don't know how I just got the answer from my teacher and she not showing me how to do it and i am asking her and she getting me upset

OpenStudy (anonymous):

lol ok i will teach u

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):



OpenStudy (anonymous):

it okay take your time

OpenStudy (anonymous):

Can you factorise equations?

OpenStudy (anonymous):

4x^2 - 4x -3 = 0 (2x-3)(2x+1)=0 2x-3=0 or 2x+1=0, so x=3/2 or x=-1/2

OpenStudy (anonymous):

thanks everyone for your help I really do accept it.

OpenStudy (anonymous):



OpenStudy (anonymous):

okay thanks guys

OpenStudy (anonymous):

can i have a medal or fan

OpenStudy (anonymous):

d